# New versions of Frobenius and integral closure of ideals

### v2

@@ 4.2. The adjusted closures in affine semigroup rings
Let $s\geq 0$ be a real number, and suppose that $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}$. Now, for any $q\geq 1$, we have that
$q\mathbf{v}+\operatorname{Exp}(\mathfrak{a}^{\lceil tq\rceil})\subset q\mathbf{v}+tqH+\sigma^{\vee}\subset q\mathcal{B}+sqH+\sigma^{\vee}\subset q\mathcal{B}+\operatorname{Exp}(\mathfrak{a}^{\lfloor sq-d\rfloor})+\sigma^{\vee}.$
Note that $(q\mathcal{B}+\operatorname{Exp}(\mathfrak{a}^{\lfloor sq-d\rfloor})+\sigma^{\vee})\cap\bar{S}=\operatorname{Exp}(I^{[q]}\mathfrak{a}^{\lfloor sq-d\rfloor}\bar{R})$.
Suppose $s=0$, so that we have $q\mathbf{v}+\operatorname{Exp}(\mathfrak{a}^{\lceil tq\rceil})\subset\operatorname{Exp}(I^{[q]}\bar{R})$, and so $x^{q\mathbf{v}}\mathfrak{a}^{\lceil tq\rceil}\subset I^{[q]}\bar{R}$. Hence, if $\mathbf{v}+tH\subset\mathcal{B}+\sigma^{\vee}$, then $x^{\mathbf{v}}\in(I\bar{R})^{F\mathfrak{a}^{t}}\subset(I\bar{R})^{*\mathfrak{a}^{t}}$.
Suppose $s=t$, and choose $x^{\mathbf{c}}\in\mathfrak{a}^{d+1}$. Now, for $q\geq 1$, we have that
$\mathbf{c}+q\mathbf{v}+\operatorname{Exp}(\mathfrak{a}^{\lceil tq\rceil})\subset\operatorname{Exp}(x^{\mathbf{c}}I^{[q]}\mathfrak{a}^{\lfloor tq-d\rfloor})\subset\operatorname{Exp}(I^{[q]}\mathfrak{a}^{\lfloor tq+1\rfloor})\subset\operatorname{Exp}(I^{[q]}\mathfrak{a}^{\lceil tq\rceil}).$
Therefore $x^{\mathbf{c}}x^{q\mathbf{v}}\mathfrak{a}^{\lceil tq\rceil}\subset I^{[q]}\mathfrak{a}^{\lceil tq\rceil}$ for all $q\geq 1$. Hence, if $\mathbf{v}+tH\subset\mathcal{B}+tH+\sigma^{\vee}$, then $x^{\mathbf{v}}\in{{}^{\mathfrak{a}^{t}}(I\bar{R})^{*}}$.
Suppose $s>t$, so that for $q\gg 0$, we have that $sq-d\geq\lceil tq\rceil$, and so $q\mathbf{v}+\operatorname{Exp}(\mathfrak{a}^{\lceil tq\rceil})\subset\operatorname{Exp}(I^{[q]}\mathfrak{a}^{\lfloor sq-d\rfloor}\bar{R})\subset\operatorname{Exp}(I^{[q]}\mathfrak{a}^{\lceil tq\rceil}\bar{R})$. Hence, if $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}$ for some $s>t$, then $x^{\mathbf{v}}\in{}^{\mathfrak{a}^{t}}(I\bar{R})^{F}$. ∎
@@
Given 4.1, if $k$ is of prime characteristic $p>0$ and $R=k[S]$ is a normal semigroup ring, then for a monomial $x^{\mathbf{v}}\in R$, we have
@@
The converse of Corollary 4.6(c) does not hold in general, that is, the $\mathfrak{a}^{t}$-tight and Frobenius closures may differ. For example, let $k$ be a field of odd prime characteristic $p$, $R=k[x,y]$, $I=\mathfrak{a}=(x^{2},y^{2})$, and $t=1$, so $tH=\mathrm{Hull}((2,0),(0,2))$. Then, $(1,1)+tH=\mathrm{Hull}((3,1),(1,3))$, and $\{(2,0),(0,2)\}+tH=\mathrm{Hull}((4,0),(0,4))$. Therefore $(1,1)+tH\subset\{(2,0),(0,2)\}+tH$, demonstrating an $\mathfrak{a}^{t}$-tight closure relation. However, $xy\notin{}^{\mathfrak{a}^{t}}I^{F}$; for any power $q$ of $p$, we have that $x^{3q}y^{q}\in(xy)^{q}\mathfrak{a}^{q}$. But, this monomial is not in $I^{[q]}\mathfrak{a}^{q}$, since all generators of the latter are of degree $4q$ but have only even exponents.
@@
Therefore, if Corollary 4.6 were to hold without the normality assumption, we would have that $y\in I^{F\mathfrak{a}^{t}}$. However, for any power $q$ of $p$ we have that
$y^{q}\mathfrak{a}^{q}=(x^{2a+3b}y^{q}\mid a+b=q)=(x^{2q}y^{q},x^{2q+1}y^{q})$
but $I^{[q]}=(x^{2q}y^{q})$. Since $x\notin R$, $x^{2q+1}y^{q}\notin I^{[q]}$. Therefore $y\notin I^{F\mathfrak{a}^{t}}$.
Independent of the characteristic of $k$, we also get similar convex geometry statements for the $\mathfrak{a}^{t}$-integral closure.
@@
Given 4.1, if $R=k[S]$ is a normal semigroup ring and $x^{\mathbf{v}}\in R$ is a monomial, then $x^{\mathbf{v}}\in\overline{I}^{\mathfrak{a}^{t}}$ if and only if $\mathbf{v}+tH\subset\operatorname{Hull}(\operatorname{Exp}I)$.
@@
As $E$ is a closed subset of $\mathbb{R}^{d}$ in the usual Euclidean topology, we may take the limit in $n$ and stay inside $E$, that is, $\mathbf{v}+tH\subset E$. Notably, this direction of the theorem does not require that $R$ be normal.
Now, suppose $\mathbf{v}+tH\subset E$, and we want to show $x^{\mathbf{v}}\in\overline{I}^{\mathfrak{a}^{t}}$. For some $n\in\mathbb{N}$, let $\mathbf{a}\in\operatorname{Exp}(\mathfrak{a}^{\lceil nt\rceil})$ be such that $x^{\mathbf{a}}$ is a nonzero monomial generator of $\mathfrak{a}^{\lceil tn\rceil}$. Then, there is a $\mathbf{w}$ in $\mathbb{N}^{r}$ such that $|\mathbf{w}|_{1}=\lceil tn\rceil$ and $\mathbf{a}=A\mathbf{w}$. Also, fix $\mathbf{c}\in\operatorname{Exp}(I^{\ell})$, for $\ell=\mu(I)-1$. We then have
@@
Notably, the term $\mathbf{v}+(t/\lceil tn\rceil)A\mathbf{w}$ is inside $\mathbf{v}+tH\subset E$ and the term $(1/n-t/\lceil tn\rceil)A\mathbf{w}$ is in $\sigma^{\vee}$ since $t/\lceil tn\rceil<t/tn=1/n$. Thus, $\mathbf{c}+n\mathbf{v}+\mathbf{a}$ is inside $\mathbf{c}+nE+\sigma^{\vee}\subset\mathbf{c}+nE\subset\mathbf{c}+\operatorname{Exp}(\overline{I^{n}})$. Finally, by the Briançon-Skoda theorem, $\overline{I^{n}}\subset I^{n-\ell}$ for all $n$. This shows $\mathbf{c}+\operatorname{Exp}(I^{n-\ell})\subset\operatorname{Exp}(I^{n})$ by our choice of $\mathbf{c}$, demonstrating that $x^{\mathbf{v}}\in\overline{I}^{\mathfrak{a}^{t}}$, as required. ∎
@@
We note that the $\mathfrak{a}^{t}$-tight and integral closure calculations above are entirely independent of characteristic. However, below we demonstrate characteristic dependence in the Frobenius closures.
4.3. Computed Examples
@@
Given 4.1, let $C=(\mathbf{c}_{1}|\mathbf{c}_{2}|\cdots|\mathbf{c}_{n})$ be the matrix whose columns are the vectors $\mathbf{c}_{i}$, which means that a vector $\mathbf{v}\in V$ is in $\sigma^{\vee}$ if and only if $C^{T}\mathbf{v}\geq\mathbf{0}$ componentwise.
@@
For any $s\geq 0$, $t>0$, and $\mathbf{v}\in S$, $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}$ if and only if for all $\mathbf{u}\in\Delta$ there exists $\mathbf{w}\in\Delta$ and $i$ such that $\mathbf{v}+tA\mathbf{u}\subset\mathbf{b}_{i}+sA\mathbf{w}+\sigma^{\vee}$, that is $C^{T}(\mathbf{v}-\mathbf{b}_{i}+A(t\mathbf{u}-s\mathbf{w}))\geq\mathbf{0}.$
4.3.1. The case $\mu(\mathfrak{a})=2$
@@
Suppose that $A=(\mathbf{a}_{1}\mid\mathbf{a}_{2})$, and hence $\Delta=\{(1-\gamma,\gamma)^{T}\mid\gamma\in[0,1]\}$. Now, for $\mathbf{u}=(1-\gamma,\gamma)^{T}\in\Delta$, $\mathbf{w}=(1-\delta,\delta)^{T}\in\Delta$, and $t,s\geq 0$, we have that
@@
Now,
@@
$\iff$ $\forall j,\;(t\gamma-s\delta)\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle\geq\langle\mathbf{c}_{j},\mathbf{b}_{i}\rangle-\langle\mathbf{c}_{j},\mathbf{v}\rangle-(t-s)\langle\mathbf{c}_{j},\mathbf{a}_{1}\rangle.$
Now for all $j$ with $\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle>0$, the condition above becomes
(4.2) $t\gamma-s\delta\geq\frac{\langle\mathbf{c}_{j},\mathbf{b}_{i}\rangle-\langle\mathbf{c}_{j},\mathbf{v}\rangle-(t-s)\langle\mathbf{c}_{j},\mathbf{a}_{1}\rangle}{\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle}.$
@@
For all $j$ with $\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle<0$, the condition above becomes
(4.3) $t\gamma-s\delta\leq\frac{\langle\mathbf{c}_{j},\mathbf{b}_{i}\rangle-\langle\mathbf{c}_{j},\mathbf{v}\rangle-(t-s)\langle\mathbf{c}_{j},\mathbf{a}_{1}\rangle}{\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle}.$
@@
For any $j$ with $\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle=0$, the condition is
(4.4) $\langle\mathbf{c}_{j},\mathbf{b}_{i}\rangle-\langle\mathbf{c}_{j},\mathbf{v}\rangle-(t-s)\langle\mathbf{c}_{j},\mathbf{a}_{1}\rangle\leq 0\quad\iff\quad\langle\mathbf{c}_{j},\mathbf{v}\rangle\geq\langle\mathbf{c}_{j},\mathbf{b}_{i}-(t-s)\mathbf{a}_{1}\rangle.$
@@
This inequaltiy does not depend on $\gamma$ or $\delta$, and so is a necessary condition on any $\mathbf{v}$ with $\mathbf{v}+tA\mathbf{u}\in\mathcal{B}+sH+\sigma^{\vee}$. For the remainder of this section we assume that this inequality holds on the vectors $\mathbf{v}$ we consider.

### v3

@@ 6.2. The adjusted closures in affine semigroup rings
Let $s\geq 0$ be a real number, and suppose that $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}$. Now, for any $e\geq 0$, we have that
$p^{e}\mathbf{v}+\operatorname{Exp}(\mathfrak{a}^{\lceil tp^{e}\rceil})\subset p^{e}\mathbf{v}+tp^{e}H+\sigma^{\vee}\subset p^{e}\mathcal{B}+sp^{e}H+\sigma^{\vee}\subset p^{e}\mathcal{B}+\operatorname{Exp}(\mathfrak{a}^{\lfloor sp^{e}-d\rfloor})+\sigma^{\vee}.$
Note that $(p^{e}\mathcal{B}+\operatorname{Exp}(\mathfrak{a}^{\lfloor sp^{e}-d\rfloor})+\sigma^{\vee})\cap\overline{S}=\operatorname{Exp}(I^{[p^{e}]}\mathfrak{a}^{\lfloor sp^{e}-d\rfloor}\overline{R})$.
Suppose $s=0$, so that we have $p^{e}\mathbf{v}+\operatorname{Exp}(\mathfrak{a}^{\lceil tp^{e}\rceil})\subset\operatorname{Exp}(I^{[p^{e}]}\overline{R})$, and so $x^{p^{e}\mathbf{v}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{[p^{e}]}\overline{R}$. Hence, if $\mathbf{v}+tH\subset\mathcal{B}+\sigma^{\vee}$, then $x^{\mathbf{v}}\in(I\overline{R})^{F\mathfrak{a}^{t}}\subset(I\overline{R})^{*\mathfrak{a}^{t}}$.
Suppose $s=t$, and choose $x^{\mathbf{c}}\in\mathfrak{a}^{d+1}$. Now, for $e\geq 0$, we have that
$\mathbf{c}+p^{e}\mathbf{v}+\operatorname{Exp}(\mathfrak{a}^{\lceil tp^{e}\rceil})\subset\operatorname{Exp}(x^{\mathbf{c}}I^{[p^{e}]}\mathfrak{a}^{\lfloor tp^{e}-d\rfloor})\subset\operatorname{Exp}(I^{[p^{e}]}\mathfrak{a}^{\lfloor tp^{e}+1\rfloor})\subset\operatorname{Exp}(I^{[p^{e}]}\mathfrak{a}^{\lceil tp^{e}\rceil}).$
Therefore $x^{\mathbf{c}}x^{p^{e}\mathbf{v}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{[p^{e}]}\mathfrak{a}^{\lceil tp^{e}\rceil}$ for all $e\geq 0$. Hence, if $\mathbf{v}+tH\subset\mathcal{B}+tH+\sigma^{\vee}$, then $x^{\mathbf{v}}\in{{}^{\mathfrak{a}^{t}}(I\overline{R})^{*}}$.
Suppose $s>t$, so that for $e\gg 0$, we have that $sp^{e}-d\geq\lceil tp^{e}\rceil$, and so $\operatorname{Exp}(I^{[p^{e}]}\mathfrak{a}^{\lfloor sp^{e}-d\rfloor}\overline{R})\subset\operatorname{Exp}(I^{[p^{e}]}\mathfrak{a}^{\lceil tp^{e}\rceil}\overline{R})$. Therefore,
$p^{e}\mathbf{v}+\operatorname{Exp}(\mathfrak{a}^{\lceil tp^{e}\rceil})\subset\operatorname{Exp}(I^{[p^{e}]}\mathfrak{a}^{\lfloor sp^{e}-d\rfloor}\overline{R})\subset\operatorname{Exp}(I^{[p^{e}]}\mathfrak{a}^{\lceil tp^{e}\rceil}\overline{R}).$
Hence, if $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}$ for some $s>t$, then $x^{\mathbf{v}}\in{}^{\mathfrak{a}^{t}}(I\overline{R})^{F}$. ∎
@@
Given 6.1, if $k$ is of prime characteristic $p>0$ and $R=k[S]$ is a normal semigroup ring, for a monomial $x^{\mathbf{v}}\in R$, we have
@@
The converse of Corollary 6.6(c) does not hold in general, that is, the Vraciu $\mathfrak{a}^{t}$-tight and Frobenius closures may differ. For example, let $k$ be a field of odd prime characteristic $p>0$, $R=k[x,y]$, $I=\mathfrak{a}=(x^{2},y^{2})$, and $t=1$, so $tH=\mathrm{Hull}((2,0),(0,2))$. Then, $(1,1)+tH=\mathrm{Hull}((3,1),(1,3))$, and $\{(2,0),(0,2)\}+tH=\mathrm{Hull}((4,0),(0,4))$. Therefore $(1,1)+tH\subset\{(2,0),(0,2)\}+tH$, demonstrating $xy\in{}^{\mathfrak{a}^{t}}I^{*}$. However, $xy\notin{}^{\mathfrak{a}^{t}}I^{F}$; for any $e\geq 0$, we have that $x^{3p^{e}}y^{p^{e}}\in(xy)^{p^{e}}\mathfrak{a}^{p^{e}}$. But, this monomial is not in $I^{[p^{e}]}\mathfrak{a}^{p^{e}}$, since all generators of the latter are of degree $4p^{e}$ but have only even exponents.
@@
Therefore, if Corollary 6.6 were to hold without the normality assumption, we would have that $y\in I^{F\mathfrak{a}^{t}}$. However, for any $e\geq 0$ we have that
$y^{p^{e}}\mathfrak{a}^{p^{e}}=(x^{2a+3b}y^{p^{e}}\mid a+b=p^{e})=(x^{2p^{e}}y^{p^{e}},x^{2p^{e}+1}y^{p^{e}})$
but $I^{[p^{e}]}=(x^{2p^{e}}y^{p^{e}})$. Since $x\notin R$, $x^{2p^{e}+1}y^{p^{e}}\notin I^{[p^{e}]}$. Therefore $y\notin I^{F\mathfrak{a}^{t}}$.
Independent of the characteristic of $k$, we also get similar convex geometry statements for the $\mathfrak{a}^{t}$-integral closure. These statements can be seen as restatements of Theorem 3.9, as the defining equations of the convex hull of the exponent set of $I$ exactly correspond to the Rees valuations of $I$ (see [BDHM24, Theorem A]).
@@
Given 6.1, if $x^{\mathbf{v}}\in\overline{I}^{\mathfrak{a}^{t}}$ then $\mathbf{v}+tH\subset\operatorname{Hull}(\operatorname{Exp}I)$, and if $k[S]$ is normal then the converse holds.
@@
As $E$ is a closed subset of $\mathbb{R}^{d}$ in the usual Euclidean topology, we may take the limit in $n$ and stay inside $E$, that is, $\mathbf{v}+tH\subset E$.
Now, suppose $k[S]$ is normal and $\mathbf{v}+tH\subset E$, and we want to show $x^{\mathbf{v}}\in\overline{I}^{\mathfrak{a}^{t}}$. For some $n\in\mathbb{N}$, let $\mathbf{a}\in\operatorname{Exp}(\mathfrak{a}^{\lceil nt\rceil})$ be such that $x^{\mathbf{a}}$ is a nonzero monomial generator of $\mathfrak{a}^{\lceil tn\rceil}$. Then, there is a $\mathbf{w}$ in $\mathbb{N}^{r}$ such that $|\mathbf{w}|_{1}=\lceil tn\rceil$ and $\mathbf{a}=A\mathbf{w}$. Also, fix $\mathbf{c}\in\operatorname{Exp}(I^{\ell})$, for $\ell=\mu(I)-1$. We then have
@@
Notably, the term $\mathbf{v}+(t/\lceil tn\rceil)A\mathbf{w}$ is inside $\mathbf{v}+tH\subset E$ and the term $(1/n-t/\lceil tn\rceil)A\mathbf{w}$ is in $\sigma^{\vee}$ since $t/\lceil tn\rceil<t/tn=1/n$. Thus,
$\mathbf{c}+n\mathbf{v}+\mathbf{a}\in\mathbf{(}c+nE+\sigma^{\vee})\cap S\subset\mathbf{c}+(nE\cap S)\subset\mathbf{c}+\operatorname{Exp}(\overline{I^{n}}).$
Finally, by the Briançon-Skoda theorem, $\overline{I^{n}}\subset I^{n-\ell}$ for all $n$. This shows $\mathbf{c}+\operatorname{Exp}(I^{n-\ell})\subset\operatorname{Exp}(I^{n})$ by our choice of $\mathbf{c}$, demonstrating that $x^{\mathbf{v}}\in\overline{I}^{\mathfrak{a}^{t}}$, as required. ∎
@@
We note that the $\mathfrak{a}^{t}$-tight and integral closure calculations above are entirely independent of characteristic, so for a fixed affine semigroup and ideal, these closure operations cannot vary as the characteristic of the underlying field varies. However, we demonstrate characteristic dependence in the $\mathfrak{a}^{t}$-Frobenius closures in the next example.
In the example below, the statements about the adjusted integral closures hold over fields of arbitrary characteristic $k$, but the statements involving the joint Hilbert-Kunz multiplicity and the adjusted tight and Frobenius closures require that $k$ is of prime characteristic $p>0$.
Example 6.10.
The joint Hilbert-Kunz multiplicity as a function of $t$ is given below.
The Hara-Yoshida $\mathfrak{a}^{t}$-integral closures are given below.
[…]
The Vraciu $\mathfrak{a}^{t}$-tight closures are given below.
(6.4) ${}^{\mathfrak{a}^{t}}I^{*}=\begin{cases}K_{1}:=I=(x^{4},x^{2}y^{3})&\text{ if }t<\frac{1}{2}\\ K_{2}:=K_{1}+(x^{5}y^{2})=(x^{4},x^{5}y,x^{2}y^{3})&\text{ if }\frac{1}{2}\leq t<1\\ K_{3}:=K_{2}+(x^{4}y,x^{4}y^{2})=(x^{4},x^{4}y,x^{4}y^{2},x^{2}y^{3})&\text{ if }t\geq 1\end{cases}$
[…]
We defer the details of this computation to the appendix below.
Appendix A Appendix: Calculating the closures in affine semigroup rings
In this section we provide the calculations behind 6.10, which demonstrate the ways in which these closures and multiplicity may be computed. We use the data of the linear algebra of the cone $\sigma^{\vee}$ and the generators of $I$ and $\mathfrak{a}$ to determine when a monomial $x^{\mathbf{v}}$ is in the various $\mathfrak{a}^{t}$-closures of $I$. We first handle the joint Hilbert-Kunz multiplicity and the $\mathfrak{a}^{t}$-integral closure, as those are simpler to calculate.
@@ Appendix A Appendix: Calculating the closures in affine semigroup rings
We continue to assume that we are in 6.1, and in particular note that $R=k[S]$ where $S\subset\mathbb{Z}^{d}$ is a semigroup, $\sigma^{\vee}=\mathrm{cone}(S)$, and $\mathbf{c}_{1},\ldots,\mathbf{c}_{n}\in\mathbb{Z}^{d}$ vectors such that $\mathbf{v}\in\sigma^{\vee}$ if and only if $\langle\mathbf{c}_{i},\mathbf{v}\rangle\geq 0$ for all $i=1,\ldots,n$. Let $C=(\mathbf{c}_{1}|\mathbf{c}_{2}|\cdots|\mathbf{c}_{n})$ be the matrix whose columns are the vectors $\mathbf{c}_{i}$, which means that a vector $\mathbf{v}$ is in $\sigma^{\vee}$ if and only if $C^{T}\mathbf{v}\geq\mathbf{0}$ componentwise.
Let $\mathcal{A}=\{\mathbf{a}_{1},\ldots,\mathbf{a}_{r}\},\mathcal{B}=\{\mathbf{b}_{1},\ldots,\mathbf{b}_{n}\}\subset S$, $\mathfrak{a}=(x^{\mathbf{a}}\mid\mathbf{a}\in\mathcal{A})$, and $I=(x^{\mathbf{b}}\mid\mathbf{b}\in\mathcal{B})$. Let $A=(\mathbf{a}_{1}|\cdots|\mathbf{a}_{r})$, and let $\Delta=\{\mathbf{u}\in\mathbb{R}_{\geq 0}^{r}\mid|\mathbf{u}|_{1}=1\}$, where $|\mathbf{u}|_{1}$ is the sum of (the absolute values of) the components of $\mathbf{u}$. Let $H=\mathrm{Hull}(\mathbf{a}_{1},\ldots,\mathbf{a}_{r})$ be the convex hull of the columns of $A$, and notice that $H=\{A\mathbf{u}\mid\mathbf{u}\in\Delta\}$.
A.1. The joint Hilbert-Kunz multiplicity
[…]
A.2. Computing the integral closures
$\overline{I}=(x^{4},x^{4}y,x^{3}y^{2},x^{2}y^{3}).$
A monomial $x^{v_{1}}y^{v_{2}}$ is in $\overline{I}^{\mathfrak{a}^{t}}$ if and only if $\mathbf{v}+tH$ is contained in the Newton polytope, where $H$ is the convex hull of $\mathbf{a}_{1}$ and $\mathbf{a}_{2}$. Since the Newton polytope is convex, it suffices to check that $\mathbf{v}+t\mathbf{a}_{1}$ and $\mathbf{v}+t\mathbf{a}_{2}$ are in the Newton polytope. Thus, $x^{\mathbf{v}}\in\overline{I}^{\mathfrak{a}^{t}}$ if and only if the following hold.
$\langle\mathbf{(}3,2),\mathbf{v}+t\mathbf{a}_{1}\rangle$ $\geq 12$ equivalently, $3v_{1}+2v_{2}+6t$ $\geq 12$
$\langle\mathbf{(}3,2),\mathbf{v}+t\mathbf{a}_{2}\rangle$ $\geq 12$ equivalently, $3v_{1}+2v_{2}+12t$ $\geq 12.$
[…]
This proves (6.2) in Example 6.10.
A.3. Computing the tight and Frobenius closures
To use Theorem 6.4 and Theorem 6.5, we will rewrite the condition $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}$ in terms of the components of the vectors.
@@ A.3. Computing the tight and Frobenius closures
For any $s\geq 0$, $t\in\mathbb{R}_{>0}$, and $\mathbf{v}\in S$, $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}$ if and only if for all $\mathbf{u}\in\Delta$ there exists $\mathbf{w}\in\Delta$ and $i$ such that $\mathbf{v}+tA\mathbf{u}\subset\mathbf{b}_{i}+sA\mathbf{w}+\sigma^{\vee}$, that is,
(A.1) $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}\quad\iff\forall\mathbf{u}\in\Delta\;\exists\mathbf{w}\in\Delta\;\exists i\;\text{s.t.}\;C^{T}(\mathbf{v}-\mathbf{b}_{i}+A(t\mathbf{u}-s\mathbf{w}))\geq\mathbf{0}$
@@
For our calculations, we will focus on the case $\mu(\mathfrak{a})=2$. Suppose that $A=(\mathbf{a}_{1}\mid\mathbf{a}_{2})$, and hence $\Delta=\{(1-\gamma,\gamma)^{T}\mid\gamma\in[0,1]\}$. Now, for $\mathbf{u}=(1-\gamma,\gamma)^{T}\in\Delta$, $\mathbf{w}=(1-\delta,\delta)^{T}\in\Delta$, and $s\geq 0$, $t\in\mathbb{R}_{>0}$, we have that
@@
Now, for a fixed $i$, $\mathbf{u}$, and $\mathbf{w}$, the containment on the right hand side of (A.1) can be rewritten as
@@
(A.2) $\iff$ $\forall j,\;(t\gamma-s\delta)\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle\geq\langle\mathbf{c}_{j},\mathbf{b}_{i}\rangle-\langle\mathbf{c}_{j},\mathbf{v}\rangle-(t-s)\langle\mathbf{c}_{j},\mathbf{a}_{1}\rangle.$
Now for all $j$ with $\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle>0$, (A.2) becomes
(A.3) $t\gamma-s\delta\geq\frac{\langle\mathbf{c}_{j},\mathbf{b}_{i}\rangle-\langle\mathbf{c}_{j},\mathbf{v}\rangle-(t-s)\langle\mathbf{c}_{j},\mathbf{a}_{1}\rangle}{\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle}.$
@@
For all $j$ with $\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle<0$, (A.2) becomes
(A.4) $t\gamma-s\delta\leq\frac{\langle\mathbf{c}_{j},\mathbf{b}_{i}\rangle-\langle\mathbf{c}_{j},\mathbf{v}\rangle-(t-s)\langle\mathbf{c}_{j},\mathbf{a}_{1}\rangle}{\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle}.$
@@
For any $j$ with $\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle=0$, (A.2) is
(A.5) $\langle\mathbf{c}_{j},\mathbf{b}_{i}\rangle-\langle\mathbf{c}_{j},\mathbf{v}\rangle-(t-s)\langle\mathbf{c}_{j},\mathbf{a}_{1}\rangle\leq 0\quad\iff\quad\langle\mathbf{c}_{j},\mathbf{v}\rangle\geq\langle\mathbf{c}_{j},\mathbf{b}_{i}-(t-s)\mathbf{a}_{1}\rangle.$
@@
Inequality (A.5) does not depend on $\gamma$ or $\delta$, and so is a necessary condition on any $\mathbf{v}$ with $\mathbf{v}+tA\mathbf{u}\in\mathbf{b}_{i}+sH+\sigma^{\vee}$. We will say that $\mathbf{v}$ and $\mathbf{b}_{i}$ satisfy (A.5) if for all $j$ with $\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle=0$, we have $\langle\mathbf{c}_{j},\mathbf{v}\rangle\geq\langle\mathbf{c}_{j},\mathbf{b}_{i}-(t-s)\mathbf{a}_{1}\rangle$.
Thus, we have that for a particular $i$, $\mathbf{u}=(1-\gamma,\gamma)^{T}$, and $\mathbf{w}=(1-\delta,\delta)^{T}$, that $\mathbf{v}+tA\mathbf{u}\in\mathbf{b}_{i}+sA\mathbf{w}+\sigma^{\vee}$ if and only if $m_{i}(\mathbf{v})\leq t\gamma-s\delta\leq M_{i}(\mathbf{v})$ and $\mathbf{v}$ satisfies (A.5).